### v1

@@ Abstract
The integration of Artificial Intelligence (AI) into Integrated Development Environments (IDEs) is reshaping software development, fundamentally altering how developers interact with their tools. This shift marks the emergence of Human-AI Experience in Integrated Development Environment (in-IDE HAX), a field that explores the evolving dynamics of Human-Computer Interaction in AI-assisted coding environments. Despite rapid adoption, research on in-IDE HAX remains fragmented which highlights the need for a unified overview of current practices, challenges, and opportunities. To provide a structured overview of existing research, we conduct a systematic literature review of 89 studies, summarizing current findings and outlining areas for further investigation.
Our findings reveal that AI-assisted coding enhances developer productivity but also introduces challenges, such as verification overhead, automation bias, and over-reliance, particularly among novice developers. Furthermore, concerns about code correctness, security, and maintainability highlight the urgent need for explainability, verification mechanisms, and adaptive user control. Although recent advances have driven the field forward, significant research gaps remain, including a lack of longitudinal studies, personalization strategies, and AI governance frameworks. This review provides a foundation for advancing in-IDE HAX research and offers guidance for responsibly integrating AI into software development.
@@ 1 INTRODUCTION
The integration of AI into IDEs has gained significant traction in recent years, driven by the promise of enhanced developer productivity, improved software quality, and more efficient workflows (Ziegler et al., 2022; Izadi et al., 2024). Recent large-scale industry surveys show that 76% of respondents are either using or planning to use AI tools in their development process (Stack Overflow, 2024). Additionally, more than 50% of respondents note such benefits as reduced time spent searching for information, faster coding and development, quicker completion of repetitive tasks, and an overall increase in productivity (JetBrains, 2024).
@@
Given the growing reliance on AI-enhanced IDEs in software engineering (e.g., Coursor  [^1] Cursor - The AI Code Editor https://www.cursor.com/), it is crucial to merge and evaluate existing evidence in the field of in-IDE HAX. Doing so will help to identify recurring themes and gaps in the literature regarding the developer experience with AI assistance.
We previously conducted a literature review on in-IDE HAX (Sergeyuk et al., 2024), provided an initial exploration of the topic and highlighted its potential. However, it did not employ a systematic methodology and covered only 36 studies, resulting in less coverage of the rapid increase in both research and industry adoption of AI in IDEs. Currently, no comprehensive review merges this growing body of work.
This review aims to address this gap by systematically synthesizing recent research on in-IDE HAX, building on the foundation of our inital review. In this study, we adopt a broader scope and deeper analysis guided by Preferred Reporting Items for Systematic reviews and Meta-Analyses (PRISMA) framework (Moher et al., 2010).
@@
Using our protocol, we identified 254 studies and thoroughly reviewed 89 that were relevant to in-IDE HAX. The goals of this review are to (a) merge the current state of knowledge regarding in-IDE HAX, (b) identify common themes and patterns in the literature, and (c) identify critical gaps that can guide future research efforts.
Our systematic literature review reveals key methodological trends, interaction paradigms, and the broader impact of AI integration in software development. Research in this domain predominantly employs qualitative (43%) and experimental (35%) approaches, often leveraging mixed methods to explore the evolving relationship between developers and AI. While the majority of studies focus on professional settings (82%), research on educational contexts (18%) remains limited, even though it is crucial for understanding long-term AI adoption and skill development.
Findings highlight two dominant AI interaction paradigms in IDEs: autocompletion -based assistance and conversational agents. Additionally, there is growing interest in hybrid models that combine real-time suggestions with interactive, context-aware support. Most studies ($\approx$53%) focus on GitHub Copilot, which primarily follows the autocompletion paradigm with chat as a supporting feature. However, other AI-powered development tools are also studied, some of which explore hybrid approaches. This demonstrates the variety of intelligent coding assistants.
The impact of AI on software development is complex. Although studies report productivity gains, they also highlight challenges such as verification overhead, automation bias, and over-reliance, particularly among novice developers. In AI-assisted code generation, concerns regarding correctness, security, and maintainability emphasize the need for explainability, verification mechanisms, and adaptive user control to ensure responsible adoption.
Our analysis highlights several areas that require further research. Many studies do not explicitly consider Software Development Lifecycle (SDLC) stages, limiting insights into AI ’s role beyond coding and debugging. Future work should examine AI assistance in earlier phases, such as requirements engineering and design, as well as later stages, such as testing and deployment. Furthermore, research on the long-term effects of AI-assisted development remains limited, particularly regarding skill retention and changes in collaboration. Personalization is another open challenge, with few studies exploring how AI can adapt to the expertise and workflows of individual developers. Finally, AI governance and ethical considerations, including bias mitigation and responsible adoption, require deeper investigation to ensure sustainable integration into software engineering.
Our study provides several contributions:
@@
  • Comprehensive synthesis of existing research. We analyze 89 studies, categorizing them based on their research goals, methodological approaches, and key findings. This synthesis reveals dominant research themes, study contexts, and trends in HAX research.
@@
  • Characterization of HAX research dimensions. We classify research efforts into three core aspects: (a) Design, focusing on how AI assistance is integrated into IDEs; (b) Impact, examining how AI affects developers’ workflows, productivity, and experience; and (c) Quality, assessing AI-generated code for security, comprehensibility, and adequacy.
@@
  • Identification of future research opportunities. By analyzing the proposed future work in the studies, we identify under-explored topics, such as personalization of AI assistance, the long-term effects of AI adoption, and ethical considerations in AI-powered software development.
@@
By addressing these aspects, our study provides a structured overview of in-IDE HAX research, informing both academic and industry efforts to design and evaluate AI-powered development tools.
@@ 2 METHOD
  • RQ 1. What are the key findings in the field of Human-AI Experience within Integrated Development Environments? To answer RQ1, we systematically examined study contexts, goals, research questions, and key findings from published reports. That allowed us to determine key research themes in the field and define methodological trends.
  • RQ 2. What aspects of software development related to HAX in IDEs have been extensively studied, and which areas remain under-explored? To address RQ 2, we defined the stages of SDLC according to the Waterfall model (Alshamrani and Bahattab, 2015). The Waterfall model was chosen for its clear delineation of development stages, which facilitated systematic categorization. We also examined three key aspects of HAX identified by a previous literature review (Sergeyuk et al., 2024). Design of HAX draws on insights from interviews, literature reviews, experiments, and other domains to understand how HAX solutions are conceptualized and implemented. Impact explores whether and how HAX affects users, and investigates the nature of those effects. Quality focuses on the ability of AI to generate secure, comprehensible, and contextually appropriate code suggestions.
@@ 2.1 Planning
  • Population: This review considered studies in the field of computer science that focus on software developers or users engaging with AI-enhanced tools within IDEs.
@@
During the systematic literature review’s planning stage, as recommended by the guidelines (Centre for Reviews and Dissemination (1995), UK), we established quality assessment criteria for studies that would be included after the initial screening phase. Our quality assessment questions were inspired by the recommendations of  Zhou et al. (2015):
@@
Each question was accompanied by the following scale: 0 — No, and not considered; 0.5 — Partially; 1 — Yes. Following recommended guidelines, we set the cutoff for further inclusion in the review at a score higher than 2. These criteria ensured a systematic evaluation of study design, methodological soundness, and the validity of findings.
@@ 2.2 Identifying and Screening
In November 2024, after finishing the planning phase, we searched the aforementioned sources, using titles, keywords, and abstracts as the basis for our queries.
We used the following base string, adapting it for each source’s specific system requirements: (”AI Assistant” OR ”AI Companion” OR ”AI-Powered Programming Tool” OR ”Code Completion Tool” OR ”Coding Assistant” OR ”Copilot” OR ”Intelligent Code Assistant” OR ”LLM-Based Coding Assistant” OR ”LLM-Powered Coding Assistant” OR ”LLM4Code” OR ”Programming Assistant” OR ”Human-AI Experience” OR ”Human-AI Co-Creation” OR ”Human-AI Collaboration” OR ”Human-AI Collaboration” OR ”Human-AI Interaction”). The full list of search string variations with corresponding libraries can be found in Supplementary materials  [^2] The dataset can be accessed at https://doi.org/10.5281/zenodo.14976737.
The initial search overlooked 28 relevant papers known to us from our expertise in the field, so we added them manually. We also included 36 papers from our previous literature survey (Sergeyuk et al., 2024).
In total, 224 papers entered the screening phase of the systematic literature review. We identified and excluded 29 duplicates appearing across multiple databases. Next, we conducted an initial screening, applying the exclusion criteria and identifying 113 papers out of the study’s scope.
@@ 2.3 Assessing Eligibility
In the next stage of the systematic literature review, we assessed the eligibility of 82 full-text articles based on their quality. As described in Subsection 2.1, we applied four evaluation criteria: Reporting, Rigor, Credibility, and Relevance. Articles were accepted only if their combined score on these criteria exceeded 2.
The quality assessment was conducted independently by the second and third authors, followed by a discussion to reconcile any discrepancies. In cases of disagreement, the first author acted as an arbitrator. This structured process ensured consensus among the the reviewers for the final quality assessment. A total of 8 papers were excluded due to insufficient quality scores.
@@ 2.4 Backward Snowballing
To complement the systematic search, we performed backward snowballing (Wohlin, 2014) on papers deemed eligible for inclusion. First, we extracted citations from these papers and screened them against the report-related criteria outlined in Subsection 2.1. References that did not meet these criteria were excluded. We next removed any papers already identified in previous steps We assessed the remaining studies against the study-level inclusion criteria described in Subsection 2.1, excluding out-of-scope papers. After two rounds of this process, no new papers were identified, resulting in 18 new in-scope papers. These studies underwent the eligibility assessment process outlined in Subsection 2.3, yielding a final total of 89 papers to be included in the review.
@@ 2.5 Data Extraction
As mentioned in Subsection 2.1, from all 89 eligible papers, we extracted data about Authors, Authors’ Affiliation, DOI, Publication Date, Venue, Goal of the Study, Research Questions, Key Findings, Future Work, Scope and SDLC Stage. In the later stages of data analysis, we added a Context column to indicate whether HAX was studied in a professional or educational setting.
@@
Data extraction yielded a comprehensive table with 89 studies in it. We provide the full table in Supplementary materials (Sergeyuk et al., 2025), and present a shortened overview of the studies in Table LABEL:tab:overview.
Table 1: List of studies included in the review
Title Authors Date
“It would work for me too”: How Online Communities Shape Software Developers’ Trust in AI-Powered Code Generation Tools Cheng et al. (2024a) 2024
“It’s Weird That it Knows What I Want”: Usability and Interactions with Copilot for Novice Programmers Prather et al. (2023) 2023
A Case Study in Engineering a Conversational Programming Assistant’s Persona Ross et al. (2023b) 2023
A Case Study on Scaffolding Exploratory Data Analysis for AI Pair Programmers Zhou and Li (2023) 2023
A Large-Scale Survey on the Usability of AI Programming Assistants: Successes and Challenges Liang et al. (2024) 2023
A Mixed Reality Approach for Innovative Pair Programming Education with a Conversational AI Virtual Avatar Manfredi et al. (2023) 2023
A Study on Developer Behaviors for Validating and Repairing LLM-Generated Code Using Eye Tracking and IDE Actions Tang et al. (2024) 2024
A Transformer-Based Approach for Smart Invocation of Automatic Code Completion de Moor et al. (2024) 2024
A User-centered Security Evaluation of Copilot Asare et al. (2024) 2024
AI Tool Use and Adoption in Software Development by Individuals and Organizations: A Grounded Theory Study Li et al. (2024) 2024
AI-Powered Chatbots and the Transformation of Work: Findings from a Case Study in Software Development and Software Engineering Súße et al. (2023) 2023
An Analysis of the Costs and Benefits of Autocomplete in IDEs Jiang and Coblenz (2024) 2024
@@ A Case Study on Scaffolding Exploratory Data Analysis for AI Pair Programmers Zhou and Li (2023) 2023
An Empirical Evaluation of GitHub Copilot’s Code Suggestions Nguyen and Nadi (2022) 2022
An Empirical Study of Code Search in Intelligent Coding Assistant: Perceptions, Expectations, and Directions Liu et al. (2024) 2024
An Empirical Study on Usage and Perceptions of LLMs in a Software Engineering Project Rasnayaka et al. (2024) 2024
An Exploratory Study on Upper-Level Computing Students’ Use of Large Language Models as Tools in a Semester-Long Project Tanay et al. (2024) 2024
Analyzing Prompt Influence on Automated Method Generation: An Empirical Study with Copilot Fagadau et al. (2024) 2024
Ansible Lightspeed: A Code Generation Service for IT Automation Sahoo et al. (2024) 2024
Anticipating User Needs: Insights from Design Fiction on Conversational Agents for Computational Thinking Penney et al. (2023) 2024
Are Prompt Engineering and TODO Comments Friends or Foes? An Evaluation on GitHub Copilot OBrien et al. (2024) 2024
Assessing the quality of GitHub copilot’s code generation Yetistiren et al. (2022) 2022
Better Together? An Evaluation of AI-Supported Code Translation Weisz et al. (2022) 2022
Can Developers Prompt? A Controlled Experiment for Code Documentation Generation Kruse et al. (2024) 2024
Case Study: Using AI-Assisted Code Generation In Mobile Teams Vasiliniuc and Groza (2023) 2023
CoLadder: Supporting Programmers with Hierarchical Code Generation in Multi-Level Abstraction Yen et al. (2023) 2023
Conversing with Copilot: Exploring Prompt Engineering for Solving CS1 Problems Using Natural Language Denny et al. (2023) 2023
Copilot Evaluation Harness: Evaluating LLM-Guided Software Programming Agarwal et al. (2024) 2024
Copilot for Xcode: Exploring AI-Assisted Programming by Prompting Cloud-based Large Language Models Tan et al. (2023) 2023
CoPrompt: Supporting Prompt Sharing and Referring in Collaborative Natural Language Programming Feng et al. (2023) 2024
Defendroid: Real-time Android code vulnerability detection via blockchain federated neural network with XAI Senanayake et al. (2024) 2024
Demystifying Practices, Challenges and Expected Features of Using GitHub Copilot Zhang et al. (2023a) 2023
Design Principles for Collaborative Generative AI Systems in Software Development Chen and Zacharias (2024) 2024

### v2

@@ Abstract
The integration of Artificial Intelligence (AI) into Integrated Development Environments (IDEs) is reshaping software development, fundamentally altering how developers interact with their tools. This shift marks the emergence of Human-AI Experience in Integrated Development Environment (in-IDE HAX), a field that explores the evolving dynamics of Human-Computer Interaction in AI-assisted coding environments. Despite rapid adoption, research on in-IDE HAX remains fragmented, which highlights the need for a unified overview of current practices, challenges, and opportunities. To provide a structured overview of existing research, we conduct a systematic literature review of 90 studies, summarizing current findings and outlining areas for further investigation.
We organize key insights from reviewed studies into three aspects: Impact, Design, and Quality of AI-based systems inside IDEs. Impact findings show that AI-assisted coding enhances developer productivity but also introduces challenges, such as verification overhead and over-reliance. Design studies show that effective interfaces surface context, provide explanations and transparency of suggestion, and support user control. Quality studies document risks in correctness, maintainability, and security. For future research, priorities include productivity studies, design of assistance, and audit of AI-generated code. The agenda calls for larger and longer evaluations, stronger audit and verification assets, broader coverage across the software life cycle, and adaptive assistance under user control.
@@ 1 INTRODUCTION
The integration of Artificial Intelligence (AI) into Integrated Development Environments (IDEs) has gained significant traction in recent years, driven by the promise of enhanced developer productivity, improved software quality, and more efficient workflows (Ziegler et al., 2022; Izadi et al., 2024). Recent large-scale industry surveys show that 76% of respondents are either using or planning to use AI tools in their development process (Stack Overflow, 2024). Additionally, more than 50% of respondents note such benefits as reduced time spent searching for information, faster coding and development, quicker completion of repetitive tasks, and an overall increase in productivity (JetBrains, 2024).
@@
Given the growing reliance on AI-enhanced IDEs in software engineering (e.g., Coursor [^1] Cursor: The AI Code Editor https://www.cursor.com/), it is crucial to merge and evaluate existing evidence in the field of in-IDE HAX. Doing so will help to identify recurring themes and gaps in the literature regarding the developer experience with AI assistance.
Recent literature reviews have provided comprehensive overviews of how Large Language Model (LLM) support software engineering tasks: summarizing model capabilities, task coverage, and evaluation techniques (He et al., 2025; Zhou et al., 2025a; Hou et al., 2024; Durrani et al., 2024). However, these reviews do not analyze the interaction experience inside developer workspaces, where the software development lifecycle takes place. Our work is, to the best of our knowledge, the first systematic review of in-IDE HAX. This lens highlights experience-level patterns and design challenges that general and capability-focused reviews on LLM in software engineering do not capture.
This work builds on our earlier survey of in-IDE HAX (Sergeyuk et al., 2024), which provided an initial exploration of the topic and highlighted its potential. In contrast, the present study is extended in both scope and methodological rigor. It applies a formal Systematic Literature Review (SLR) process, following the Preferred Reporting Items for Systematic reviews and Meta-Analyses (PRISMA) framework (Moher et al., 2010). Compared to the 36 papers in the initial study, which did not follow a systematic protocol, it includes 90 papers identified through a multi-step procedure that combines database search, backward snowballing, and expert-informed additions. Beyond expanding the dataset, this review introduces several analytical contributions. It not only organizes findings along three key research dimensions: Design, Impact, and Quality. It also provides a detailed analysis of the study context, empirical methods, and proposed future directions. These contributions position this study not as a follow-up review, but as a foundational reference for cumulative and reproducible progress in the research field of in-IDE HAX.
@@
Using our protocol, we identified 257 studies and thoroughly reviewed 90 that were relevant to in-IDE HAX. The goals of this review are to (a) identify common themes and patterns in the literature, (b) merge the current state of knowledge regarding in-IDE HAX, and (c) identify critical gaps that can guide future research efforts. We focus on how the presence of an AI assistance changes what happens in the very place where code is written, inspected, or tested, rather than how AI helps in software engineering broadly.
Our review reveals the focus of in-IDE HAX research activity on professional software development settings, examining the code implementation stage of the Software Development Lifecycle (SDLC), often using GitHub Copilot as the primary example. Other lifecycle stages, such as requirements, testing, and deployment, receive limited attention. Educational contexts and perspectives of non-users or people who have stopped using AI assistance are also rarely studied.
According to our review, key findings in the field span three core dimensions: Impact (effects on developers, tasks, and workflows), Design (how AI is integrated into the IDE and interacted with), and Quality (properties of AI outputs such as correctness, security, and readability). Impact-oriented studies (74/90) report productivity improvements, especially among experienced users, but also note increased time spent on result verification and concerns about over-reliance. Studies in the Design dimension (28 out of 90 papers) examine autocompletion and conversational systems, with some exploring emerging hybrid approaches. Research highlights the influence of prompt structure and interface properties such as context awareness, transparency, and user control on experience with AI tooling. In the Quality dimension (19/90), studies evaluate correctness, maintainability, and security of generated code, often calling for improved validation and diagnostic support.
Future research directions identified in the literature focus on several key areas, including productivity factors, audit mechanisms for AI-generated code, and the design of AI assistance tools. There is a strong emphasis on personalization and transparency, with a growing interest in exploring the long-term adoption of AI tools. Methodologically, studies suggest the need for larger-scale evaluations, comparative research across user groups and contexts, and longitudinal studies to track the evolution of AI’s impact over time. Despite significant progress, many research directions remain underexplored, particularly in areas such as AI governance, user control, and proactivity.
In summary, while substantial progress has been made in understanding in-IDE HAX, key gaps remain, particularly in exploring non-user perspectives and underrepresented stages of the software development lifecycle. Future research should focus on broadening the contexts studied, incorporating longitudinal designs, and refining methodologies to provide a more comprehensive understanding of AI’s long-term impact and its role across diverse user groups.
Our study offers the following contributions:
@@
  • Comprehensive synthesis of existing research. We analyze 90 studies, categorizing them based on their research goals, methodological approaches, and key findings. This synthesis reveals dominant research themes, study contexts, and trends in HAX research.
@@
  • Characterization of HAX research dimensions. We classify research efforts into three core aspects: (a) Impact, examining how AI affects developers’ workflows, productivity, and experience; (b) Design, focusing on how AI assistance is integrated into IDEs; and (c) Quality, assessing AI-generated code for security, comprehensibility, and adequacy.
@@
  • Identification of future research opportunities. By analyzing the proposed future work in the studies, we identify under-explored topics, such as AI governance, user control, and proactivity.
@@
By addressing these aspects, we provide a structured overview of in-IDE HAX research, informing both academic and industry efforts to design and evaluate AI-powered development tools.
@@ 2 METHOD
  • RQ 1. What aspects of software development related to HAX within IDEs have been extensively studied, and which areas remain under-explored? To address RQ1, we defined the stages of SDLC according to the Waterfall model (Alshamrani and Bahattab, 2015). The Waterfall model was chosen for its clear delineation of development stages, which facilitated systematic categorization. We also examined study contexts and three key aspects of HAX identified by a previous literature review (Sergeyuk et al., 2024).
  • RQ 2. What are the key findings in the field of HAX in IDEs? To answer RQ2, we systematically examined goals, research questions, and key findings from published reports. That allowed us to determine key research themes in the field and define methodological trends.
@@
To address RQ1 and RQ2, we applied an in-IDE HAX taxonomy (Sergeyuk et al., 2024) developed through a staged qualitative procedure. Two authors independently conducted open coding of the corpus to propose candidate categories, followed by consensus meetings to merge or split codes, assign names, and draft a codebook with definitions and examples. The codebook was piloted on a subset, refined, and then used to code the full dataset by two authors, with disagreements adjudicated by a third. Categories with low support or conceptual overlap were consolidated to improve parsimony while preserving coverage. The resulting structure comprises three distinct dimensions: Impact (effects on developers, tasks, and workflows), Design (how AI is integrated into and interacted with in the IDE), and Quality (properties of AI outputs such as correctness, security, and readability).
@@ 2.1 Planning
  • Population: We considered studies in the field of computer science that focus on software developers or users engaging with AI-enhanced tools within IDEs.
@@
IDE in this review denotes any workspace that simultaneously offers (a) code editing, (b) immediate execution or preview, and (c) contextual AI feedback in the same window. Traditional desktop IDEs (IntelliJ, VS Code) and notebook-style environments (Jupyter, Databricks) satisfy all three criteria; plug-ins that embed a full editor and run code inline also qualify. Tools that push AI suggestions outside the primary editing surface (e.g., web explainers, offline static analysers) are excluded.
During the planning stage, as recommended by the guidelines (Centre for Reviews and Dissemination (1995), UK), we established quality assessment criteria for studies that would be included after the initial screening phase. Our quality assessment questions were inspired by the recommendations of Zhou et al. (2015):
@@
Each question was accompanied by the following scale: 0 — No, and not considered; 0.5 — Partially; 1 — Yes. Following recommended guidelines, we set the cutoff for further inclusion in the review at a score higher than 2. These criteria ensured a systematic evaluation of the research’s aims, study design, methodological soundness, value and validity of findings.
@@ 2.2 Identifying and Screening
In November 2024, after finishing the planning phase, we searched the aforementioned sources, using titles, keywords, and abstracts as the basis for our queries. We optimized for high recall at the search stage and enforced precision during manual screening.
The search string has a two-part structure: a core query covering terminology related to AI assistants and Human-AI Experience (Score), optionally combined with a query specifying IDE-related terms (Side).
  • Score: “AI Assistant” OR “AI Companion” OR “AI-Powered Programming Tool” OR “Code Completion Tool” OR “Coding Assistant” OR “Copilot” OR “Intelligent Code Assistant” OR “LLM-Based Coding Assistant” OR “LLM-Powered Coding Assistant” OR “LLM4Code” OR “Programming Assistant” OR “Human-AI Experience” OR “Human-AI Co-Creation” OR “Human-AI Collaboration” OR “Human-AI Interaction”
  • Side: “Integrated Development Environment” OR “Code Editor” OR “Coding Environment” OR “Development Environment” OR “Programming Environment”
The full query was used in IEEE Xplore, Web of Science, ScienceDirect, Scopus, Springer Link, and arXiv. In ACM Digital Library and DBLP, we used only Score due to limited filtering capabilities and higher baseline relevance. We did not include “developer” as a required keyword, since it tended to retrieve general developer studies and exclude relevant work focused on tools or interfaces that met our criteria but were not indexed using that term. All search string variations and their application per library are available in the Supplementary materials (Sergeyuk et al., 2025).
We complemented our database search with two additional inclusion strategies. First, we included 35 studies from our prior literature survey (Sergeyuk et al., 2024), which, although not systematic, provided coverage of early work in this fast-moving field. Second, we manually added 30 relevant studies known to the authors through domain expertise. These were identified as frequently cited or influential in recent research but were missed due to inconsistent metadata, terminology drift, or indexing limitations in digital libraries. This expert-based supplementation is recommended in SLR methodology (Wohlin, 2014; Kitchenham et al., 2009) to reduce omission of contextually critical work. All manually included studies were subjected to the same quality assessment and eligibility criteria as search-derived papers to ensure methodological consistency.
All retrieved records were manually screened against our in-IDE HAX inclusion criterion. This step preserved recall while restoring precision. In total, 223 papers entered the screening phase of the systematic literature review. We identified and excluded 30 duplicates appearing across multiple databases. Next, we conducted an initial screening, applying the exclusion criteria and identifying 114 papers out of the scope of the current SLR.
@@ 2.3 Assessing Eligibility
In the next stage of the systematic literature review, we assessed the eligibility of 79 full-text articles based on their quality. As described in Subsection 2.1, we applied four evaluation criteria: Reporting, Rigor, Credibility, and Relevance. Articles were accepted only if their combined score on these criteria exceeded 2.
The quality assessment was conducted independently by the second and third authors, followed by a discussion to reconcile any discrepancies. In cases of disagreement, the first author acted as an arbitrator. This structured process ensured consensus among the reviewers for the final quality assessment. A total of 8 papers were excluded due to insufficient quality scores lower or equal to 2, dictated by insufficient relevance and rigor. Excluded papers were accessed by authors as only partially or not sufficient in terms of clarity of the statement of the aims of the research, its design, findings related to the aims of the research, and value for research or practice.
@@ 2.4 Backward Snowballing
To complement the systematic search, we performed backward snowballing (Wohlin, 2014) on papers deemed eligible for inclusion. First, we extracted citations from these papers and screened them against the report-related criteria outlined in Subsection 2.1. References that did not meet these criteria were excluded. We next removed any papers already identified in previous steps We assessed the remaining studies against the study-level inclusion criteria described in Subsection 2.1, excluding out-of-scope papers. After the first round of this process, 18 new in-scope papers were identified. A second round of snowballing did not yield any additional relevant papers. These 18 studies underwent the eligibility assessment process outlined in Subsection 2.3.
During the revision phase, we re-verified the included studies. This process led to the removal of 2 papers (one out-of-scope, one duplicate) and the addition of 3 previously overlooked studies, ensuring that the final set reflects both the inclusion criteria and recent relevant literature and yielding a final total of 90 papers to be included in the review.
Note that eligibility was determined by the date of first public availability. We included studies first available between January 2022 and November 2024, including preprints on arXiv. When a study later appeared in an archival venue in 2025, we cite the archival version with its DOI and venue year. As a result, some entries list 2025 in the reference list while remaining within the time window of our review based on the preprint date.
@@ 2.5 Data Extraction
As mentioned in Subsection 2.1, from all 90 eligible papers, we extracted data about Authors, Authors’ Affiliation, DOI, Publication Date, Venue, Goal of the Study, Research Questions, Key Findings, Future Work with its topic and methodology, Scope, SDLC Stage. In the later stages of data analysis, we added a Context column to indicate whether HAX was studied in a professional or educational setting. Moreover, we classified studies based on the types of tasks investigated in the work.
@@
Data extraction yielded a comprehensive table with 90 studies in it. We provide the full table in Supplementary materials (Sergeyuk et al., 2025), and present a shortened overview of the studies in Table LABEL:tab:overview.
Table 1: List of studies included in the review. Dating note. Inclusion used first public availability between January 2022 and November 2024. We cite the archival version with its DOI and venue year. Some entries show 2025 because an earlier preprint falls within this window.
ID Title Authors
@@
1 An Empirical Evaluation of GitHub Copilot’s Code Suggestions Nguyen and Nadi (2022)
2 Asleep at the Keyboard? Assessing the Security of GitHub Copilot’s Code Contributions. Pearce et al. (2022)
3 Assessing the quality of GitHub copilot’s code generation Yetistiren et al. (2022)
4 Better Together? An Evaluation of AI-Supported Code Translation Weisz et al. (2022)